OpenStudy (anonymous):

Find the missing angle measure in triangleABC. angleA = 94°, angleB = 37°, angleC = ____° A. 94 B. 37 C. 149 D. 49

OpenStudy (anonymous):

D

OpenStudy (***[isuru]***):

the sum of interior angles of a triangle = 180 so ... angle A + angle B + angle C = 180 94 + 37 + angle C = 180 angle C = 180 - 137 = 49 got it :) ?

OpenStudy (anonymous):

I love this stuff. :D

OpenStudy (***[isuru]***):

lol XD

OpenStudy (anonymous):

:D